OpenStudy (anonymous):

Which exponential equation is equivalent to the logarithmic equation below? log 478= a A.10^a=478 B.478^10=a C.478^a=10 D.a^10=478

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log(478)=a }\) (so the presumed base is 10) is this correct?

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \displaystyle \log_{ \color{green}{\rm a}}\color{blue}{\rm b}=\color{red}{\rm c} ~~~~\Longrightarrow~~~~\color{green}{\rm a}^{\color{red}{\rm c}}=\color{blue}{\rm b}}\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \displaystyle \log_{ \color{green}{\rm 10}}\color{blue}{\rm 478}=\color{red}{\rm a} ~~~~\Longrightarrow~~~~?}\)